The sum of Mark's and his Mom's age is 48. His Mom's age is three times his age. What is the age of each?
Answer by JulietG(1812)

K + M = 48 where K is Mark and M is Mom
M = 3K
Substitute the value of M from the second equation into the first
K + M = 48
K + (3K) = 48
4K = 48
Divide each side by 4
K = 12
If Mark is 12, then his mother is 36 (3x his age).